How many grams of sodium bicarbonate, NaHCO3, are needed to make 200.0 ml of a 0.250 M solution?


1
Expert's answer
2022-06-07T06:06:01-0400

n(NaHCO3)=c×V=0.2×0.25=0.05 mol

M(NaHCO3)=23+1+12+48=84 g/mol

m(NaHCO3)=n×M=0.05×84=4.2 g